OpenStudy (ilikestudying140898):

What is the 4th term of 112 square root 2 a4 -112 square root 2 a4 112 square root 2 a5 -112 square root 2 a5

OpenStudy (ilikestudying140898):

Binomial Theorem

OpenStudy (anonymous):

-112

OpenStudy (ilikestudying140898):

thanks

OpenStudy (ilikestudying140898):

@abideinhim square root a4 or a5

OpenStudy (mathmale):

If you know Pascal's Triangle and its applications, use that tool to find the coefficients.

OpenStudy (anonymous):

\[a^8-8 \sqrt{2} a^7+56 a^6-112 \sqrt{2} a^5+280 a^4-224 \sqrt{2} a^3+224 a^2-64 \sqrt{2} a+16 \]\[-112 \sqrt{2} a^5 \]

imqwerty (imqwerty):

the general term \(T_{r-1}\) of any binomial expansion say \((a+b)^n\) is given as-> \(\large T_{r-1}=~^nC_r(a)^{n-r}(b)^r\) you wanna find \(4^{th}\) term or we can say \(T_4\) so \(T_{r-1}=T_4\) so r=5 now just find out \(a,~b,~n\) and put them in the general term equation along with r to get the answer :)

OpenStudy (ilikestudying140898):

the answer is D